OpenStudy (anonymous):

You have the numbers 1–24 written on slips of paper. If you choose one slip at random, what is the probability that you will not select a number that is divisible by 6? 1/6 5/6 1/3 2/3

OpenStudy (anonymous):

@Jacob,*

OpenStudy (anonymous):

try doing the same thing as last time,

OpenStudy (anonymous):

number of favorable outcomes divided by the total number of outcomes

OpenStudy (anonymous):

Numbers which are divisible by 6 = 6,12,18,24 P (number divisible by 6) = 4/24 P (number not divisible by 6) = \[1-\frac{ 4 }{ 24 }\]

OpenStudy (anonymous):

\[1-\frac{ 4 }{ 24 }\] isent one of the choices though?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

no it's 1 minus 4 over 24

OpenStudy (anonymous):

you have to do the equation, but it should e pretty easy

OpenStudy (anonymous):

\[1 -\frac{ 4 }{ 24 } = \frac{ 24-4 }{ 24 } \] \[\frac{ 24-4 }{ 24 } = \frac{ 20 }{ 24 } =\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

think of 1 as \[\frac{ 24 }{ 24 }\]

OpenStudy (anonymous):

thank you im sorry i got coufused

OpenStudy (anonymous):

np